Find the P-Value for the indicated Hypothesis test: A medical school claims that more than 28% of its students plan to go into general practice. It is found that among a random sample of 130 of the school's students, 32% of them plan to go into general practice. Find the P-value for a test of the school's claim
Answer by stanbon(75887)

Find the P-Value for the indicated Hypothesis test:
A medical school claims that more than 28% of its students plan to go into general practice. It is found that among a random sample of 130 of the school's students, 32% of them plan to go into general practice.
Find the P-value for a test of the school's claim
---
z(0.32) = (0.32-0.28)/sqrt[0.28*0.72/130] = 1.0157
----
p-value = P(z > 1.0157) = 0.1549
===============
Cheers,
Stan H.
===============